Passes through (10,2), perpendicular to y= 4x + 6
Answer by fractalier(6550)

The line perpendicular to y = 4x + 6 has a slope of (-1/4).
Plug all of the info into y = mx + b to find b...
2 = (-1/4)(10) + b
2 = -5/2 + b
b = 9/2
so that the line we want is
y = (-1/4)x + 9/2